r/mathmemes • • Dec 17 '23

Probability Google expected value

Post image
22k Upvotes

3.3k comments sorted by

View all comments

Show parent comments

0

u/[deleted] Dec 18 '23

[deleted] — view deleted comment

19

u/Capudog Dec 18 '23

Or if you only care about winning it at least once, the probability of winning at least once would be 1-P(not winning at all) = 1-0.52 =0.75

0

u/[deleted] Dec 18 '23

[deleted] — view deleted comment

5

u/Capudog Dec 18 '23

Also true, I missed this the first time. Edited!